### v1

# HA-RDet: Hybrid Anchor Rotation Detector for Oriented Object Detection

Abstract
Oriented object detection in aerial images poses a significant challenge due to their varying sizes and orientations. Current state-of-the-art detectors typically rely on either two-stage or one-stage approaches, often employing Anchor-based strategies, which can result in computationally expensive operations due to the redundant number of generated anchors during training. In contrast, Anchor-free mechanisms offer faster processing but suffer from a reduction in the number of training samples, potentially impacting detection accuracy. To address these limitations, we propose the Hybrid-Anchor Rotation Detector (HA-RDet), which combines the advantages of both anchor-based and anchor-free schemes for oriented object detection. By utilizing only one preset anchor for each location on the feature maps and refining these anchors with our Orientation-Aware Convolution technique, HA-RDet achieves competitive accuracies, including 75.41 mAP on DOTA-v1, 65.3 mAP on DIOR-R, and 90.2 mAP on HRSC2016, against current anchor-based state-of-the-art methods, while significantly reducing computational resources. Source code: https://github.com/PhucNDA/HA-RDet
1 Introduction
Object detection in aerial images (ODAI) stands as a pivotal challenge within the domains of computer vision and pattern recognition, boasting numerous practical applications [19, 18, 20, 6, 28, 14]. Furthermore, there exists a pressing need within the research community to explore advanced technologies capable of autonomously analyzing vast-scale remote sensing images. However, unlike scenes captured from horizontal perspectives, aerial images obtained from an overhead view often feature densely populated objects appearing in various scales, shapes, and arbitrary orientations. Such complexities pose formidable challenges for models, especially those relying on Horizontal Bounding Boxes (HBB) to accurately represent object instances in the image. To address this, Oriented Bounding Boxes (OBB) are utilized to offer a more precise depiction of object locations by incorporating additional directional information. Notably, several significant challenges emerge in the realm of detecting objects in aerial images, including:
  • Large aspect ratio: Objects in aerial images often have irregular shapes, resulting in high aspect ratios, such as bridges, ships, harbors, etc.
  • Scale variations: Varying ground sampling distance (GSD) of sensors leads to scale variation among objects captured by different sensors in the same scene [23].
  • Dense arrangement: Aerial images usually contain multiple objects distributed densely, such as ships in a harbor or vehicles in a parking lot, which can pose a big challenge for object detection algorithms as they need to accurately identify and distinguish object instances.
  • Arbitrary orientations: Objects in aerial images can appear in various orientations, requiring object detection models to have precise orientation estimation capabilities.
Existing oriented detectors [5, 26, 9, 25, 3] typically rely on proposal-driven frameworks utilizing a Region Proposal Network (RPN), which employs numerous anchors with varying scales and aspect ratios to generate Regions of Interest (RoIs). However, horizontal RoIs often cover multiple instances, causing ambiguity, while rotated RoIs improve recall but incur high computational costs. Hence, developing an efficient region proposal network capable of generating high-quality proposals is crucial to overcome accuracy and computational limitations in current state-of-the-art oriented detectors. We identify two key reasons for the lack of efficiency in region proposal-based oriented detectors: (1) Most oriented object detection methods adopt Anchor-based or Anchor-free training approaches, each with its own challenges and trade-offs in terms of performance and computational efficiency. Anchor-based methods achieve good performance but are sensitive to oriented anchor hyperparameters, while Anchor-free methods accelerate training speed but suffer from decreased detection performance due to a shortage of level features. (2) The conventional convolution technique, designed with fixed receptive fields and aligned with axes, is ill-suited for objects with arbitrary orientations and varying scales in aerial images. Predefined anchor representations often fail to fully capture the shape and orientation of objects, leading to misalignment between fixed convolution features and object orientations, thus hindering effective object localization and classification in oriented object detection models.
This study introduces a novel method called Hybrid-Anchor Rotation Detector (HA-RDet) for detecting oriented objects, which combines the strengths of Anchor-based and Anchor-free mechanisms through a Hybrid anchor training approach, employing a single anchor per location on the feature map. In the anchor-based component, we introduce a new Orientation-aware Convolution (O-AwareConv) technique that enhances proposal quality by adapting to object shape and orientation. Through extensive experimentation on established oriented object detection datasets and comparison with previous detectors, we demonstrate the effectiveness of our approach. Our method addresses challenges such as diminishing training samples and over-encoding information by reducing anchor hyperparameters initially and refining them later to generate high-quality proposals. Additionally, we design a lightweight transformation network in the final R-CNN head to extract rotated proposals from horizontal ones. Unlike prior studies, HA-RDet exhibits superior efficiency in both training and testing, requiring fewer computational resources while achieving promising accuracy and speeding up detection. This work aims to offer an elegant and efficient region proposal-based oriented detector that bridges the gap between anchor-free and anchor-based approaches, serving as a robust baseline for future investigations.
Our contributions are summarized as follows:
  • Presenting the Hybrid-Anchor Rotation Detector that innovatively combines anchor-based and anchor-free techniques for oriented object detection
  • Introducing the novel O-AwareConv for enhanced orientation-sensitive feature extraction
  • Designing a simple yet effective light-weight network for orientation transformation
  • Extensive experiments are conducted on challenging oriented object detection datasets to demonstrate the effectiveness of our proposed method.
2 Related works
With the surge of machine learning, particularly deep learning, object detection has witnessed significant advancements, dividing into two main categories: two-stage detectors and one-stage detectors. Two-stage detectors begin by generating a sparse set of Regions of Interest (RoIs) in the first stage, followed by RoI-wise bounding box regression and object classification in the second stage. In contrast, one-stage detectors directly detect objects without the need for RoI generation. Furthermore, two fundamental training strategies, Anchor-free and Anchor-based methods, are crucial for object detection. While Anchor-free methods excel in rapid object detection, they suffer from a scarcity of positive training samples, resulting in performance degradation. Conversely, Anchor-based methods generate numerous anchors, facilitating an abundant supply of positive training samples and achieving excellent performance, albeit at a higher computational cost. Our HA-RDet integrates both training strategies, utilizing only one preset anchor for each feature map’s location, thereby significantly reducing resource requirements while maintaining asymptotically state-of-the-art performance.
2.1 Anchor-Based Detectors
Most methods in object detection rely on predefined generated anchors, where multiple preset anchors with varying scales and ratios are assigned to each specific location on a feature map. For instance, ROI Transformer generates preset horizontal anchors and then applies RRoI Leaner and RRoI Wrapping for robust feature extraction. Oriented R-CNN proposed by Xie et al. [25] adopts an anchor-based approach with an Oriented Region Proposal Network, along with the Midpoint-offset representation. This representation encodes horizontal anchors into six distinguishable learnable parameters, ensuring a reasonable transformation from horizontal generated anchors to oriented proposals. Similarly, Faster R-CNN [20] introduces the Region Proposal Network with nine multi-scale heuristically tuned anchors for every location on a feature map. These densely generated anchors are assigned either positive or negative labels based on the Intersection over Union (IoU) with a ground-truth box. Additionally, SSD [15] enhances detector efficiency by directly predicting object location offsets and corresponding categories. While anchor-based methods are well-studied and yield remarkable performance by generating a large number of anchors with different ratios and scales for each location, they require significant computational resources for training and inference, making them impractical for oriented object detection in aerial images datasets.
2.2 Anchor-Free Detectors
The anchor-free mechanism differs from anchor-based strategies by utilizing only one key point for each location on a feature map, implying that each point on the feature map corresponds to a single anchor with a specific scale and ratio. $S^{2}$-Net [8] introduced the single-shot alignment network featuring the Feature Alignment Module and Oriented Detection Module, which ensure feature alignment using the Alignment Convolution specifically designed for oriented feature extraction. YOLO [19] divides the image into an $S\times S$ grid, with each grid cell responsible for determining the presence of an object’s center. CornerNet [11] predicts the final bounding box using a pair of key points (top-left and bottom-right) and undergoes a complex post-processing phase to combine key points of the same object. FCOS [21] assigns all locations within an object’s ground truth as positive, incorporating a ”centerness” branch to suppress low-quality bounding boxes, thereby enhancing overall performance. Despite being faster, anchor-free methods may experience a performance decline compared to previous anchor-based methods.
3 Proposed Hybrid-Anchor Rotation Detector
Figure 1: Illustration of the architecture of the proposed HA-RDet. At first, an aerial image is passed through the FPN backbone to extract deep features. Then, the extracted features are fed through the Hybrid-Anchor RPN to produce horizontal anchors via an Anchor-free assigner strategy on rectangularized oriented ground truths. The anchors are refined with the novel Orientation-aware Convolution and the later Anchor-based head to produce high-quality horizontal proposals. After that, the lightweight proposal transformation network learns to produce oriented proposals from predicted RoI Align processed horizontal proposals. Finally, these oriented proposals are refined through oriented bounding box heads for classification and bounding box regression.
Most oriented object detectors follow the anchor-based strategy for predicting the heavily encoded oriented bounding boxes. They often generate multiple anchors of different ratios and scales to allocate many instances of various shapes and sizes greedily. This approach requires a large number of anchors, leading to high computational requirements. On the other hand, the anchor-free mechanism uses only a single anchor per location on a feature map, but it suffers from accuracy deterioration due to the lack of positive training samples. Many studies improve the performance of either method via post-refining, new assigner strategies, new anchor definitions, etc. Two strategies standstill and need to be fully exploited in oriented object detection.
In this section, we reason our proposals and explain how the strategy is applied and operated appropriately under different anchor encoder-decoder fashion.
3.1 Hybrid-Anchor RPN
Anchor encoder-decoder. We first investigate HA-RDet adopting RetinaNet [14], initially designed for generic horizontal object detection but adapted for oriented object detection, where anchors are encoded as 5D $(x,y,w,h,\theta)$. Utilizing the Anchor-free assigner, the Anchor-free head aims to train the oriented anchor generator from feature maps extracted via regular 2D convolution (2DConv). Subsequently, the Anchor-based head employs a 5D encoder-decoder strategy to regress offset values $(\delta x,\delta y,\delta w,\delta h,\delta\theta)$ for the final oriented proposals. However, this experimental approach proves ineffective due to two primary reasons: (1) Insufficient positive samples hinder the model from acquiring adequate information, resulting in insufficient oriented proposals; (2) Excessive information encoded within the oriented anchors in the Anchor-free head prevents the anchor generator from converging.
Our proposed approach (depicted in Fig. 1) tackles these challenges by introducing the Hybrid-Anchor RPN, which simplifies anchor encoding from 5D $(x,y,w,h,\theta)$ to 4D $(x,y,w,h)$ while ensuring consistency between the two heads. We achieve this by utilizing the Anchor-free assigner strategy on rectangularized oriented ground truths for horizontal anchors, significantly increasing the number of positive samples. This augmentation aids the model in learning sufficient information for further refinement. Subsequently, the resulting horizontal anchors undergo processing in the Anchor-based stage, producing high-quality horizontal proposals for downstream tasks. This design guarantees an adequate number of positive samples for training and gradually enhances the quality of the final proposals. Detailed insights into the algorithmic implementation of the Hybrid-Anchor RPN are provided in Algorithm 1.
Anchor-Free head. We use one preset anchor with a single scale and ratio for each location on a feature map. In the Anchor-free head, the normal RetinaNet [14] architecture used as a backbone encodes and decodes the horizontal anchors $(a_{x},a_{y},a_{w},a_{h})$ with the rectangularized ground truth target $(t_{x},t_{y},t_{w},t_{h})$ as the following Eq. 1:
$\begin{cases}\delta_{x}=(t_{x}-a_{x})/a_{w},&\delta_{y}=(t_{y}-a_{y})/a_{h}\\ \delta_{w}=\log(t_{w}/a_{w}),&\delta_{h}=\log(t_{h}/a_{h})\\ a^{\prime}_{x}=\widehat{\delta_{x}}a_{w}+a_{x},&a^{\prime}_{y}=\widehat{\delta_{y}}a_{h}+a_{y}\\ a^{\prime}_{w}=a^{\prime}_{w}\exp(\widehat{\delta_{w}}),&a^{\prime}_{h}=a^{\prime}_{h}\exp(\widehat{\delta_{h}})\\ \end{cases}$ (1)
where $\delta=(\delta_{x},\delta_{y},\delta_{w},\delta_{h})$ and $\widehat{\delta}=(\widehat{\delta_{x}},\widehat{\delta_{y}},\widehat{\delta_{w}},\widehat{\delta_{h}})$ are predicted encoded value and target encoded value, respectively. Having regressed $\widehat{\delta}$, we can easily decode the proposal $(a^{\prime}_{x},a^{\prime}_{y},a^{\prime}_{w},a^{\prime}_{h})$. Having the classification branch omitted, the Anchor-free head aims to minimize the bounding box loss function $L_{AF}$:
$L_{AF}(\delta,\widehat{\delta})=w^{AF}[-\ln(\frac{Intersection(\delta,\widehat{\delta})}{Union(\delta,\widehat{\delta})})]$ (2)
where $w^{AF}$ is the loss weight used for fine-tuning the model.
Anchor-Based head. The Anchor-based head, defined by Eq. 1, receives anchors generated from the preceding stage along with their extracted features from Orientation-aware Convolution. Its objective is to generate refined, high-quality proposals. The Anchor-based bounding box loss function $L_{AB}$ is derived from Eq. 2, albeit with a different weight $w^{AB}$. However, different from Anchor-free, the classification branch with binary cross-entropy loss in this stage is used. The overall loss of our Region Proposal Network is the sum of the Anchor-free loss $L_{AF}$ and the Anchor-based loss $L_{AB}$, denoted as $L_{RPN}$.
$L_{RPN}=L_{AF}+L_{AB}$ (3)
Orientation-aware Convolution. Object detection in aerial images presents unique challenges compared to standard perspective images due to densely populated instances, diverse aspects, and varying scales. Models tailored for aerial object detection require precise localization and robust feature extraction.
Our introduced Orientation-Aware Convolution (O-AwareConv) is utilized in anchor-based heads, aiming to achieve global feature alignment by adaptively considering the object’s ground truth orientation $\theta_{GT}$. The proposed method involves designing a convolution that extensively calculates the shape offset value for each horizontal anchor in conjunction with the corresponding ground-truth orientation offset value. During training, a multi-level horizontal anchor list is generated, and O-AwareConv computes the shape offset field $(x,y,w,h)$ for each anchor to adjust the aligned features’ sampling locations. Additionally, orientation offset values $R(\theta_{GT})^{T}$ are adapted based on the corresponding ground-truth orientation $\theta_{GT}$, where $T$ denotes matrix transposition. This process allows convolutional features at a given location to be transformed into oriented features based on the specific anchor and ground-truth target, as depicted in Fig. 2. In summary, for each location $p$ on a feature map $X$, O-AwareConv considers the corresponding anchor $a^{*}=(x,y,w,h)$, stride $S$, kernel weight $w$, and kernel size $k$, which is represented as Eq. 4:
$\begin{cases}Y(p)=\sum_{r\in Kernels;o\in\Omega}w(r)\cdot X(p+r+o)\vspace{0.75em}\\ \Omega=\frac{1}{S}(\textit{{(x,y)}}+\frac{1}{k}\textit{{(w,h)}}\cdot r\cdot R({\theta_{GT}})^{T})-p-r\vspace{0.75em}\\ R(\alpha)=\begin{bmatrix}cos{\alpha}&-sin{\alpha}\\ sin{\alpha}&cos{\alpha}\end{bmatrix}\end{cases}$ (4)
In our approach, each anchor corresponds positively to only one specific ground truth, whereas each ground truth may be associated with multiple anchors. Leveraging every training sample, or training anchor, our method retains negative anchors in anchor-free heads as a set of potential predefined candidates, with the expectation of gradual improvement during the refining stage. During testing, where corresponding ground truth is unavailable, the orientation offset field is set to zero when calculating the final features offset.
Figure 2: The shape offsets are computed via anchor size; ground-truth orientation is adopted to perform orientation offsets calculation. Shape offsets and orientation offsets produce the final features used for training the Anchor-based refinement stage. The final computed features are orientation-awareness, surging HA-RDet performance significantly.
Figure 3: Left: Detection results on DOTA. Right: Detection results on HRSC2016
1
Input:
  • X: Multi-level features
  • gt: Ground truth
  • rgt: Rectangularized ground truth
Output:
  • Region proposals
Anchor-free stage
 1. A = anchor_generator(X)
 2. target = assigner(A, rgt)[’index’]
 3. X = 2Dconv(X)
 4. bbox_pred = regression(X)
 5. A = decode(A, bbox_pred)
 6. loss_af = box_loss(A, rgt[target])
Anchor-based stage
 7. target = assigner(A, rgt)[’index’]
 8. offset = shape_offset(A)
 9. if training phase:
10.    or_offset = or_offset(A, gt[target])
11.    offset = dot(offset, or_offset)
12. X = OAconv(X, offset)
13. bbox_pred = regression(X)
14. bbox_cls = classification(X)
15. A = decode(A, bbox_pred)
16. loss_rgs = box_loss(A, rgt[target])
17. loss_cls = class_loss(bbox_cls, label[target])
18. loss_ab = loss_rgs + loss_cls
Return
19. loss_rpn = loss_af + loss_ab
20. proposals = refinement(A, bbox_cls)
APPENDIX
  • target: Target ground-truth
  • loss_af: Anchor-free loss
  • loss_ab: Anchor-based loss
  • decode(*): Bounding box decoder
  • dot(*): Dot product mathematical operation
  • assigner(*): Anchor assigner strategy
  • OAconv(*): Orientation-aware Convolution
  • shape_offset(*): Get Shape offset
  • or_offset(*): Get Orientation offset
  • refinement(*): Box refinement via NMS
Algorithm 1 Hybrid-Anchor RPN
Compared to regular 2D convolution, which assumes objects in an image align with the axes, our approach utilizes Deformable Convolution (DeformConv) [2] with augmented offsets. However, DeformConv may inaccurately sample from incorrect locations, particularly with weak supervision for densely distributed objects. To address this limitation, we introduce the Orientation-Aware Convolution during the training phase for every generated horizontal anchor, ensuring global convergence of the Hybrid-Anchor RPN. Unlike DeformConv and 2DConv, offset field in O-AwareConv is inferred from the actual shape and orientation of objects. Fig. 2 demonstrates that our O-AwareConv accurately extracts features. In inference, global features are extracted effectively, enhancing anchor prediction accuracy without the need for the orientation feature channel.
3.2 R-CNN oriented bounding box heads
Having regressed well-qualified horizontal proposals from the Hybrid-Anchor RPN, we implemented a lightweight multi-layer neural network to transform the horizontal to oriented region proposals. RoIAlign [10] with bilinear interpolation is adopted to our baseline to extract the region of interest features. We then use 5D offset value representation $(\delta x,\delta y,\delta w,\delta h,\delta\theta)$ to encode the oriented proposals from horizontal ones and perform simple regression using two fully-connected layers.
For the oriented bounding box head, we adopt the Rotated RoIAlign operation [25] to extract rotation-sensitivity features from each oriented proposal.
4 Experiments
4.1 Datasets
DOTA-v1.0 [24] is a large-scale dataset comprising high-quality aerial images for detecting objects, containing $2,806$ high-defined images with resolutions $4000\times 4000$. It classifies annotated instances into $15$ common categories: plane (PL), baseball diamond (BD), bridge (BR), ground track field (GTF), small vehicle (SV), large vehicle (LV), ship (SH), tennis court (TC), basketball court (BC), storage tank (ST), soccer ball field (SBF), roundabout (RA), harbor (HA), swimming pool (SP) and helicopter (HC). During the training and testing process, the images are divided into cropped patches of size $1024\times 1024$ pixels with a stride of $824$ pixels, including random flipping is applied as a form of data augmentation.
DIOR-R [1] is an extended version of the DIOR dataset [12] that shares the same 23,463 optimal remote sensing images but is labeled with oriented bounding boxes. The images are $800\times 800$ pixels and vary in spatial resolution. Each object in the image is categorized into one of $20$ classes. The training procedure for DIOR-R follows the same one implemented on the DOTA-v1.0 dataset.
HRSC2016 [16] dataset is a collection of 1061 high-resolution aerial images with oriented bounding box annotations for ship detection. The image sizes range from $300\times 300$ to $1500\times 900$. This dataset encompasses over $25$ categories of ships with a large diversity of shapes, scales, ratios, and orientations annotated. During the training process, the images are resized to $800\times 512$ pixels while preserving their original aspect ratio. Additionally, random flipping is applied to augment the training data.
Model Backbones PL BD BR GTF SV LV SH TC BC ST SBF RA HA SP HC mAP
One-stage
$R^{3}$et [27] ResNet50+FPN 89.3 75.31 45.41 69.24 75.54 72.89 79.29 90.89 81.02 83.26 58.82 63.15 63.41 62.21 37.41 69.8
RSDet [17] ResNet50+FPN 89.3 82.7 47.7 63.9 66.8 62 67.3 90.8 85.3 82.4 62.3 62.4 65.7 68.6 64.6 70.8
$S^{2}$-Net [8] ResNet50+FPN 89.3 80.49 50.42 73.23 78.42 77.4 86.8 90.89 85.66 84.24 62.16 65.93 66.66 67.76 53.56 74.19
Two-stage
RoI Transformer [5] ResNet101+FPN 88.64 78.52 43.44 75.92 68.81 73.68 83.59 90.74 77.27 81.46 58.39 53.54 62.83 58.93 47.67 69.56
Gliding Vertex [26] ResNet50+FPN 89.21 75.92 51.31 69.57 78.12 75.66 86.87 90.9 85.4 84.78 53.36 66.66 66.31 69.99 54.4 73.23
Oriented R-CNN [25] ResNet50+FPN 89.35 81.41 52.6 75.02 79.03 82.41 87.82 90.9 86.4 85.3 63.36 65.7 68.28 70.48 57.23 75.69
ReDet [9] ReResNet50+ReFPN 88.79 82.64 53.97 74.00 78.13 84.06 88.04 90.89 87.78 85.75 61.76 60.39 75.96 68.07 63.59 76.25
Anchor-free
Rotated RepPoints [28] ResNet50+FPN 83.36 63.71 36.27 51.58 71.06 50.35 72.42 90.1 70.22 81.98 47.46 59.5 50.65 55.51 53.07 59.15
CFA [7] ResNet101+FPN 89.26 81.72 51.81 67.17 79.99 78.25 84.46 90.77 83.4 85.54 54.86 67.75 73.04 70.24 64.96 75.05
AOPG [1] ResNet50+FPN 89.27 83.49 52.50 69.97 73.51 82.31 87.95 90.89 87.64 84.71 60.01 66.12 74.19 68.30 57.80 75.24
Ours
HA-RDet ResNet50+FPN 88.98 83.34 55.02 72.04 78.25 81.43 86.94 90.91 85.46 85.48 61.33 61.57 74.50 68.97 56.86 75.408
HA-RDet ResNet101+FPN 89.12 84.95 54.81 69.33 77.67 82.91 87.54 90.87 86.06 86.03 64.19 61.90 74.25 70.99 65.43 76.02
HA-RDet ResNeXt101_DCNv2+FPN 89.31 85.01 57.03 68.94 78.14 83.35 87.61 90.91 87.32 86.67 65.14 63.05 76.35 72.29 63.98 77.012
Table 1: Comparison of SOTA results on the oriented object detection DOTA 1.0 task.
Model R3Det [27] AO2-DETR [3] Gliding Vertex [26] $S^{2}$-Net [8] ReDet [9] AOPG [1] Oriented RepPoints [13] Oriented R-CNN [25] HA-RDet
#Anchors 21 - 20 1 20 1 9 20 1
mAP 87.14 88.12 88.20 90.14 90.2 90.34 90.38 90.4 90.20 / 95.32
Table 2: Results on HRSC2016 with #Anchors represents the quantity of anchors generated per location on a feature map.
Model R3Det [27] CFA [7] $S^{2}$-Net [8] Rotated FCOS [21] RoI Transformer [5] AOPG [1] Oriented R-CNN [25] HA-RDet
mAP 57.2 60.8 61.7 62.4 63.9 64.4 65.1 65.3
Table 3: Experimental comparison of SOTA methods on DIOR-R, trained using ResNet50+FPN
Model (anchor scale) Backbones PL BD BR GTF SV LV SH TC BC ST SBF RA HA SP HC mAP
HA-RDet (4) ResNet50 + FPN 88.98 83.34 55.02 72.04 78.25 81.43 86.94 90.91 85.46 85.48 61.33 61.57 74.50 68.97 56.86 75.408
HA-RDet (6) ResNet50 + FPN 89.01 81.84 54.34 73.72 72.59 77.52 86.88 90.9 87.66 85.75 61.55 61.91 75.69 66.54 59.72 75.032
HA-RDet (8) ResNet50 + FPN 88.97 80.79 53.18 70.64 70.96 76.49 78.98 90.9 85.57 84.58 59.29 60.83 75.91 66.49 53.28 73.122
Table 4: Models trained on DOTA with an anchor’s scale larger than 6 perform better on large-instance classes such as GTF, HA, HC, PL, etc. However, small-instance classes, such as SV, LV, SH, etc., have higher performance on models with the anchor’s scale of less than 6.
4.2 Implementation details
We utilize two backbone architectures: ResNet50+FPN and ResNet101+FPN, as well as ResNext101-FPN-DCNv2, all pretrained on ImageNet [4]. Our optimization setup includes the SGD optimizer with an initial learning rate of 0.005 and a decay factor of 10 for each decay step. The weights $w^{AF}$ and $w^{AB}$ are both set to 7.0. We set the momentum to 0.9 and the weight decay to 0.0001. The anchor’s scale and ratio is set to 4.0 and 1.0, respectively. For the anchor assigner strategy, we employ the Ratio-based strategy (positive: 0.3; negative: 0.1) [22] for the Anchor-Free head and the Max-IoU strategy (positive: 0.7; negative: 0.3) for the Anchor-based head. Both assigning strategies are followed by random balanced sampling of 256 positive and 256 negative samples. We train HA-RDet for 12 epochs on the DOTA and DIOR-R datasets and 36 epochs on the HRSC2016 dataset. Training is conducted using a single A5000 GPU with a batch size of 2, while inference is performed on a single A5000 GPU.
4.3 Comparisons with State-of-the-Arts
The experimental results presented in Tab. 1 demonstrate the performance of HA-RDet compared to previous detectors on the DOTA-v1 dataset. HA-RDet achieves state-of-the-art asymptotic results with 75.41 mAP and 76.02 mAP using the regular ResNet50 and ResNet101 backbones, respectively. The highest result of 77.01 mAP was particularly impressive with ResNeXt101_DCNv2, surpassing most existing one-stage and anchor-free methods. By utilizing O-Aware convolution and aligning features accordingly, our model significantly improves the accuracy of detecting irregularly shaped objects such as bridges, storage tanks, and harbors, efficiently addressing high aspect ratio challenges.
On the HRSC2016 dataset, as shown in Tab. 2, HA-RDet achieves competitive results of 90.20 mAP and 95.32 mAP on the VOC2007 and VOC2012 metrics, respectively. These results outperform other training methods that use multiple anchors and demonstrate competitiveness against the $S^{2}$-Net, which follows the same training scheme of only using one anchor per location.
Tab. 3 presents the outcomes obtained from the DIOR-R dataset, where our novel Hybrid-Anchor training schemes approach demonstrates superior performance. Using the backbone ResNet50, our method achieves the highest accuracy at 65.3 mAP, surpassing the well-known two-stage Oriented R-CNN by 0.2 mAP, the one-stage $S^{2}$-Net by 3.6 mAP, and the Anchor-free Oriented Proposal Generator (AOPG) by 0.9 mAP.
We conduct further experiments to compare our proposed Hybrid-Anchor training scheme with the well-known anchor-free $S^{2}$-Net and anchor-based Oriented R-CNN methods in Tab. 5. We evaluate these methods beyond the mean Average Precision (mAP) measure, such as inference speed (FPS), computational resources (VRAM) for storing as well as processing the intermediate data during the training process, and the number of model parameters. The $S^{2}$-Net reached the detection result of 74.19 mAP with a fast inference speed of 15.5 FPS; it consumed 4.6 GB of VRAM due to its relatively low number of parameters. Meanwhile, Oriented R-CNN, with 20 anchors in the training stage, attained a slightly better accuracy than $S^{2}$-Net but with a trade-off regarding inference speed (13.5 FPS) and higher computational resource requirements (14.2 GB of VRAM). Our HA-RDet follows the Hybrid-Anchor training approach and generates only one anchor per location, similar to $S^{2}$-Net. It achieved better detection accuracy (75.41 mAP) while maintaining a stable inference speed of 12.1 FPS. Compared to Oriented R-CNN, our model has a higher parameter requirement for processing but consumes approximately half the amount of VRAM (6.8 GB) due to the significant reduction in the number of generated anchors. Oriented R-CNN utilized dense generated anchors and a complex encoding strategy; therefore, its accuracy and inference speed is slightly higher than our model, but it comes at the cost of significantly higher computational resource requirements. Some visualization of HA-RDet are demonstrated in Fig. 3
$S^{2}$-Net [8] Oriented R-CNN [25] HA-RDet (Ours)
#Anchors 1 20 1
mAP 74.19 75.69 75.41
FPS 15.5 13.5 12.1
VRAM (GB) 4.6 14.2 6.8
#params 38,599,976 41,139,754 55,715,449
Table 5: Evaluation results of HA-RDet with $S^{2}$-Net and Oriented R-CNN on the ResNet50+FPN on DOTA
Convolution method mAP
Standard Convolution 74.1
Deformable Convolution [2] 74.5 ($\uparrow$ 0.4)
Orientation-aware Convolution (Ours) 75.41 ($\uparrow$ 1.3)
Table 6: Comparison of Orientation-aware Convolution accuracy with other convolutions on the DOTA-v1 dataset.
4.4 Ablation studies
Effectiveness of O-AwareConv. As mentioned in section 3.1, we introduced a groundbreaking approach called O-AwareConv that enhances the awareness of ground-truth orientation. Tab. 6 compares O-AwareConv against alternative methods under the same configurations to demonstrate its effectiveness in oriented object detection. Our HA-RDet employs Standard Convolution and reaches only 74.1 mAP because the fixed receptive field fails to extract features adequately. On the other hand, Deformable Convolution addresses feature misalignment, yet its performance degrades when detecting densely packed and arbitrarily oriented objects in aerial images. O-AwareConv significantly enhances the performance of our proposed detector, surpassing Deformable Convolution by 0.91 mAP and outperforming standard Convolution by 1.31 mAP.
Effectiveness of Hybrid-Anchor mechanism. To evaluate the effectiveness of the Anchor-free head, Anchor-based head, and O-AwareConv in our RPN, we experiment with different HA-RDet settings where omitting certain branches, shown in Tab. 7. When we excluded the Anchor-free head and relied solely on the anchor-based mechanism, the accuracy slightly decreased to 73.153 mAP compared to using the entire component on DOTA-v1. The role of the Anchor-free head is to provide initial proposals not constrained by predefined anchor boxes. By removing this module, we argue that the detector loses its adaptability to various object sizes and aspect ratios, which can lead to a decrease in performance.
Module Our Hybrid-Anchor RPN
Anchor-free head -
Anchor-based head -
O-AwareConv - -
mAP 75.408 0.038 73.153
Table 7: Experimental evaluation of our Hybrid-Anchor RPN by taking out a couple of modules, implemented on ResNet50+FPN on DOTA-v1.
On the other hand, taking out the Anchor-based head indicates a terrible accuracy of 0.038 mAP. The Anchor-based head is responsible for refining the sparsely initial proposals and generating high-quality region proposals. Without this component, the detector lacks the refinement process, leading to imprecise and less reliable region proposals. Furthermore, without the Anchor-based head, the detector heavily relies on the initial proposals generated by the Anchor-free head, which may not be as effective in handling objects with different scales, aspect ratios, or complex backgrounds.
These configurations neglect the use of O-Aware Convolution, which bridges the Anchor-free and Anchor-based heads in our Hybrid-Anchor RPN. By disregarding one of these modules, the detector loses the refinement process, accurately predicting bounding box regression, and the ability to handle object shape and orientation information, resulting in a significant decline in performance.
Study on anchor scale. We use one preset anchor with one scale and ratio for each location on a feature map following the anchor-free mechanism Tab. 4. Then, we extend the training longer, and the diminishing of samples happens in models having anchor scale larger than 6 due to the overwhelming number of small-sized instances of the datasets. Finally, we decided to choose the anchor scale of 4 with ratio of 1.0 as the representative for our HA-RDet baseline because of its stability, easy convergence, and suitability for high-quality aerial images.
5 Conclusions
We propose the Hybrid-Anchor training scheme detector called HA-RDet with the novel O-AwareConv to tackle the issues of limited positive samples and over-encoding information in oriented object detection. We extensively experiment on three challenging aerial image datasets, achieving competitive accuracy and inference time against state-of-the-art methods with negligible computational resources.